# Less is More: Local Intrinsic Dimensions of Contextual Language Models

### v1

@@ 4.3 Local Dimensions Detect Exhaustion of Training Capabilities
To demonstrate the application of our local dimension estimates in this setting, we compute it for the sequence-tagging-based dialogue state tracking model TripPy-R (Heck et al.,, 2022) trained on the MultiWOZ2.1 dataset (Eric et al.,, 2020). In dialogue state tracking, the task is to predict the user’s intent from the natural language input utterances, and keep track of the user’s goal described throughout the conversation by updating the dialogue state (Young et al.,, 2010). The contextualized hidden states for computing our measure are taken from the last layer of the unified encoder component, which has the RoBERTa architecture. The actual state tracking is performed based on classifier outputs derived from this encoder’s output. The encoder is fine-tuned during the TripPy-R training via the loss signal derived from predicting the current dialogue turn’s state update. We train the models for 20 epochs with Adam, with a linear learning rate schedule up to $5\cdot 10^{-5}$ that starts with one warm-up epoch.
[…]
@@
Figure 3: Development of Trippy-R performance measures (model loss in green; joint goal accuracy in orange) compared with mean local dimension estimates (dark blue) evaluated on the training, validation, and test split of the MultiWOZ dataset. We show the mean and standard deviation of the measures evaluated at the end of each epoch over six different model training seeds.
@@ Results
Figure 3 shows the development of two model performance measures: One is the differentiable model loss, which is used for backpropagation. The other is the non-differentiable joint-goal accuracy (JGA), which is used to judge the downstream performance of the state tracking model.
Note that the mean local dimension estimates behave similarly when comparing the training, validation, and test splits. This is in contrast to the loss and JGA: On the training set, the loss (green) is monotonically decreasing, while the JGA (orange) is increasing. On the other hand, the loss reaches a minimum on the validation set after $7500$ batches, long before the JGA converges. In this case, the differentiable model validation loss would give the wrong impression that the model has finished learning generalizable representations. However, from the dimension estimate, we conclude that the model has not converged at that point. The mean local dimension is still decreasing, and one should continue training. Notably, we can make this observation on the training data alone, and would not need a separate labeled validation set when relying upon our estimates.
In the other direction, we can see that later in the training, after roughly 25 0002500025\,00025 000 batches, the stabilization of the local dimension estimates coincides with the convergence of the model performance regarding JGA on the validation and test sets. This effect of stabilization of the local estimates predicting that the training capabilities have been exhausted can be seen even more clearly for longer training runs (50 epochs) in Figure 16 in Section C.4.
@@ 4.4 Local Dimensions Detect Overfitting
[…]
@@
Figure 4: Development of emotion recognition model performance measures (loss in green; weighted F1 in orange; macro F1 in red) compared with mean local dimension estimates (dark blue) evaluated on the training, validation, and test split of the EmoWOZ dataset. We show the mean and standard deviation of the measures evaluated at the end of each epoch over four different training seeds.
@@ Results
Figure 4 depicts the model’s performance measures plotted against the mean local estimates. After the first training epoch, the model’s mean local dimension drops notably to $\approx 7.25$ from the base model’s initial dimension of $\approx 9.94$. Over the subsequent eight training epochs, this dimension increases to a value of about 8. Note that this behavior of the mean local estimates can be observed on all three data splits. As expected, the training loss monotonically decreases, while both weighted and macro F1-scores increase on the train set.
However, the minimum of the local dimension, followed by an increase, suggests that the model has already found an efficient representation after a single training epoch. We previously observed a connection between rising dimension and memorization in the grokking experiments in Section 4.2. This suspicion is confirmed when considering the performance measures on the validation set in Figure 4(b). After the first epoch, the validation loss increases strongly over the following epochs; thus, the model is clearly overfitting on the training examples.
@@ 4.5 Computational Resources
Fine-tuning of the RoBERTa-base and GPT-2-medium models (in Section C.1) on one of our selected datasets can be performed efficiently on a single NVIDIA V100 or GTX1080TI-12GB GPU within a few hours. Additionally, computing the embeddings for a single layer of these models requires only forward passes, which takes approximately 10 minutes on the same hardware.
@@ 4.6 Limitations
While versatile, our framework also has several limitations. One drawback is its comparatively high computational complexity. Hence, to get a reliable dimension estimate, we need to build a sufficiently large sub-sample of the latent space by performing forward passes on the input data and subsequently computing the neighborhoods and estimates. Being independent of the model training procedure, these computations can be performed in parallel to any potential fine-tuning, thus not slowing down the training process. Further work might involve making the local estimate computation more efficient and developing local descriptors that share the stability properties but are computationally faster to evaluate. An intrinsic limitation arises from the assumptions of our dimension estimates. The TwoNN estimate, which we locally apply in the neighborhoods, is known to accurately reflect the true local dimension only under strong assumptions on the local point sampling distribution. (The point sample should come from a Poisson point process on a subspace of uniform dimension.) However, this is only of secondary importance for our method, as we are more interested in changes of dimension than in absolute values. In addition, our framework is general and permits the use of other estimation methods. Finally, while the TwoNN method is non-parametric, the precise absolute value of the mean TwoNN estimate depends on the hyperparameter choices in selecting the latent space sub-sample and neighborhood size. This implies that our mean of local dimensions is of a “relative” nature, because we cannot directly compare the dimensions’ values between different model architectures. For fine-tuning runs, one should instead look at the relative change compared to the starting value.
@@ 5 Conclusion
We introduce a novel geometric perspective on LLM training dynamics by measuring the local intrinsic dimension of contextual token embeddings. While the latent spaces of contextual language models exhibit regional variation in dimensionality, the mean local dimension provides a stable, interpretable summary across dataset splits. Across diverse tasks, a sustained drop in mean local dimension reliably predicts improved generalization, offering a practical, unsupervised diagnostic signal that, among other things, enables the prediction of grokking. Such a marked reduction in intrinsic dimension in supervised downstream tasks highlights their strong compressive effect on the latent space. On the other hand, rising dimensions can point to a tradeoff between generalization and specialization, typically observed during overfitting. Our approach enables monitoring without supervision by labels, which is particularly valuable in low-resource settings where validation labels are limited or unavailable. Thus, our findings highlight the utility of geometric descriptors for monitoring and interpreting LLM behavior beyond traditional label-based metrics.
@@
To the best of our knowledge, we are the first to investigate local dimensional shifts and their implications in training and fine-tuning contextual language models for different downstream tasks. The findings of this work enhance the ongoing discussion regarding the interpretability, adaptability, and generalizability of LLMs. Our work provides a foundation for designing better model architectures and developing interventions that utilize the insight that lower intrinsic dimensions benefit machine learning problems.
The locality of the dimension estimates promises further applications beyond the scope of the current work. One immediate avenue for investigation is to see to what extent the dimension can be used to detect on which data the model has been trained. Another important direction would be to find connections between the underlying meaning that the tokens carry and their corresponding dimension estimates.
@@ Appendix A Sensitivity Analysis
  • In Section A.1, we show that with increasing sample and neighborhood sizes, the TwoNN estimates tend to converge in our setting, and already a small portion of samples suffices for reasonable reliability.
  • In Section A.2 we show that the mean local estimates are stable under applying noise to the embedding vectors.
  • In Section A.3, we show that for masked language models, both masked and regular token embeddings lead to similar mean estimates.
All embeddings in the following analyses are computed using the regular token embeddings from the last layer of the RoBERTa base model.
@@ A.1 Dependence on hyperparameter choices
Note that in our sub-sampling, we shuffle the dataset split using a given seed, and then truncate to a beginning segment of the specified size. Thus, when comparing different data sequence sub-sampling sizes for a fixed seed, the smaller sequence sub-samples are subsets of the larger ones. In all the experiments in this section, the size of token sub-sample is set to $N=$ 60 0006000060\,00060 000, and the local neighborhood size parameter $L=128$. Note that for the smallest sequence sub-sample size $M=$ $2000$, not enough non-padding tokens are necessarily available to reach a token sample size of $N=$ 60 0006000060\,00060 000, but with more sequences, we get this common token space size.
We take sequence sub-samples of each given size with 5 different seeds and select a random token sub-sample for each of these, on which the local TwoNN estimates are based. For the results, see Figure 5 and Figure 6.
For a given sequence sample size, the average of the mean local estimates, calculated across different sampling seeds, remains consistent between the various dataset splits of the same dataset. This observation indicates that the sampling process is robust to differences in how the data is divided. Furthermore, the estimates stabilize as the sequence sample size increases, suggesting convergence of the local mean estimates with larger samples. Based on these observations, we set $M$ to the values $7000$ or 10 0001000010\,00010 000 for the natural language-based tasks in Sections 4.1, 4.3 and 4.4, because it represents a practical compromise between computation effort and stability of the value at that scale. This value provides sufficient sampling density to capture the underlying data distribution comparably across datasets while maintaining computational efficiency.
@@
Figure 5: Boxplots of the mean local TwoNN estimates for different sequence sample sizes $M$ ranging from $2000$ to 16 0001600016\,00016 000. We compute the mean estimates for 5 different sequence sub-sampling seeds for the three splits of the MultiWOZ dataset. Here and in all subsequent boxplots, the average of the mean estimates is depicted as green triangles, their median is the orange line; outliers are shown as circles.
@@
Here, we want to investigate the influence of the token sub-sample size on the local TwoNN estimate. We conduct our analysis on the MultiWOZ dataset. Since here we are interested in the effect of the second sub-sampling process, we want to fix the first sampling step for the sequences: We take the sub-sample of the first $M=$ 10 0001000010\,00010 000 sequences from the MultiWOZ training, validation, and test split. From the resulting token embedding spaces, we then sample 100 000100000100\,000100 000 points (ignoring padding and end-of-sequence (EOS) tokens). Finally, we compute the local estimates based on the first $N=$ 10 0001000010\,00010 000, 20 0002000020\,00020 000, …, 90 0009000090\,00090 000, 100 000100000100\,000100 000 points in the given sub-sample. Note that we are left with slightly fewer points in the local estimation step for the largest sub-sample size because of our de-duplication step in algorithm 1. This means that in a given token sub-sample, the beginning segment of the first 10 0001000010\,00010 000 points is the same even when increasing the sample size, and we compare the mean of the local estimates of this beginning segment.
[…]
(a) MultiWOZ training data.
(b) MultiWOZ validation data.
(c) MultiWOZ test data.
@@
Having fixed sample sizes $M$ and $N$, Figure 8 shows the distribution of TwoNN dimension estimates across varying locality scale parameters, represented by the number of neighbors $L$. A key observation is the relative stability of the estimates as the number of neighbors increases, with smaller locality scales (fewer neighbors) yielding more diverse estimate distributions. This indicates that local properties of the data significantly influence dimension estimates.
@@
(a) MultiWOZ validation data.
(b) Reddit validation data.
Figure 8: Distribution of the local token-wise TwoNN estimates over different locality scales $L$ (number of neighbors). Here shown are the mean and quartiles of the resulting measure varying over the different tokens in a fixed subsample. The more global the estimates, the smaller the standard deviation in the resulting distribution.

### v2

@@ 4.3 Local Dimensions Detect Exhaustion of Training Capabilities
To demonstrate the application of our local dimension estimates in this setting, we compute them for the sequence-tagging-based dialogue state tracking model TripPy-R (Heck et al.,, 2022) trained on the MultiWOZ2.1 dataset (Eric et al.,, 2020). In dialogue state tracking, the task is to predict the user’s intent from the natural language input utterances, and keep track of the user’s goal described throughout the conversation by updating the dialogue state (Young et al.,, 2010). The contextualized hidden states for computing our measure are taken from the last layer of the unified encoder component, which has the RoBERTa architecture. The actual state tracking is performed based on classifier outputs derived from this encoder’s output. The encoder is fine-tuned during the TripPy-R training via the loss signal derived from predicting the current dialogue turn’s state update. We train the models for 20 epochs with Adam, with a linear learning rate schedule up to $5\cdot 10^{-5}$ that starts with one warm-up epoch.
The input data for the local dimension estimates is formatted in the same way as it is used during training and inference. A single input sequence consists of a dialogue turn, followed by the dialogue history, where the different sequence components are separated via special tokens. For the local dimension estimates, we sample M=7000M=\$7000\$ sequences from the training, validation, and test split, with a token subsample size N=60  [^0] N=\$60\,000\$, and neighborhood size L=128L=\$128\$.
@@
Figure 3: Development of TripPy-R performance measures (model loss in green; joint goal accuracy in orange) compared with mean local dimension estimates (blue) evaluated on the training, validation, and test split of the MultiWOZ dataset. We show the mean and standard deviation of the measures evaluated at the end of each epoch over six different model training seeds.
@@ Results
Figure˜3 shows the development of two model performance measures: One is the differentiable model loss, which is used for backpropagation. The other is the non-differentiable joint-goal accuracy (JGA), which is used to judge the downstream performance of the state tracking model.
Note that the mean local dimension estimates behave similarly when comparing the training, validation, and test splits. This is in contrast to the loss and JGA: On the training set, the loss (green) is monotonically decreasing, while the JGA (orange) is increasing. On the other hand, the loss reaches a minimum on the validation set after $7500$ batches, long before the JGA converges. In this case, the differentiable model validation loss would give the wrong impression that the model has finished learning generalizable representations. However, from the dimension estimate, we conclude that the model has not converged at that point. The mean local dimension is still decreasing, and one should continue training. Notably, we can make this observation on the training data alone and would not need a separate labeled validation set when relying upon our estimates.
Quantitatively, taken over the whole time series progression and step-wise averaged over six seeds, the mean local dimension exhibits a strong negative Spearman rank correlation with the JGA of 0.982-0.982/0.958-0.958/0.905-0.905 for the training/validation/test split, respectively. This confirms that progressive compression of the embedding space directly reflects improvements in generalizable task performance. Correlation between the mean local dimension and the loss is positive $(0.982)$ on the training set, and negative (0.612-0.612/0.734-0.734) on the validation and test set, mirroring the fact that the loss does not align consistently with JGA performance over the different splits.
In another direction, we can see that later in the training, after roughly 25 00025\,000 batches, the stabilization of the local dimension estimates coincides with the convergence of the model performance regarding JGA on the validation and test sets. This effect of stabilization of the local estimates, indicating that the training capabilities have been exhausted, can be seen even more clearly for longer training runs (50 epochs) in Figure˜17 in Section˜C.4.
@@ 4.4 Local Dimensions Detect Overfitting
We use the EmoWOZ dataset (Feng et al.,, 2022), with the task of classifying dialogue utterances into one of seven emotion classes. Our experiments employ the ERToD model (Feng et al.,, 2023) in its baseline configuration, a linear classifier built upon a bert-base-uncased (Devlin et al.,, 2019) model feature extractor. We intentionally decide on the setup without dialogue history, data augmentation, or auxiliary predictions, which are modifications proposed by the authors to solve the problem of classifying rare emotions. The model is trained for 8 epochs using a linear learning rate schedule with warmup for AdamW and peak learning rate of $2\cdot 10^{-5}$. At the end of each training epoch, local dimension estimates are computed from hidden states extracted at the last BERT-encoder layer. Thus, this is the latent space that the model has tuned to aid its subsequent classification. We select representative subsets of M=7000M=\$7000\$ sequences from the training, validation, and test split, with a token subsample size N=60  [^0] N=\$60\,000\$, and neighborhood size L=128L=\$128\$. Model performance is evaluated using cross-entropy loss and weighted F1 and macro F1 classification scores. To better capture effects on minority classes, the majority “neutral” class is excluded from the F1-scores.
@@
Figure 4: Development of emotion recognition model performance measures (loss in green; weighted F1 in orange; macro F1 in red) compared with mean local dimension estimates (blue) evaluated on the training, validation, and test split of the EmoWOZ dataset. We show the mean and standard deviation of the measures evaluated at the end of each epoch over four different training seeds.
@@ Results
Figure˜4 depicts the model’s performance measures plotted against the mean local estimates. After the first training epoch, the model’s mean local dimension drops notably to $\approx 7.25$ from the base model’s initial dimension of $\approx 9.94$. Over the subsequent eight training epochs, this dimension increases to a value of about 8. Note that this behavior of the mean local estimates can be observed on all three data splits. As expected, the training loss monotonically decreases, while both weighted and macro F1-scores increase on the train set.
However, the minimum of the local dimension, followed by an increase, suggests that the model has already found an efficient representation after a single training epoch. We previously observed a connection between rising dimension and memorization in the grokking experiments in Section˜4.2. This suspicion is confirmed when considering the performance measures on the validation set in Figure˜4(b). After the first epoch, the validation loss increases strongly over the following epochs; thus, the model is clearly overfitting on the training examples.
@@
Note that in general, the evaluation criteria are task-specific: In contrast to the observations in Section˜4.3, where the validation loss is not a reliable stopping criterion when aiming for the target metric JGA, here, the validation loss and weighted F1 are tightly aligned. Hence, the loss serves as a valid stopping signal here. The local dimension measures support this point, since they are negatively correlated with the training loss (Spearman rank 0.952-0.952), but positively correlated with the validation ($0.952$) and test loss ($0.976$).
@@ 4.5 Computational Resources
Fine-tuning of the RoBERTa-base and GPT-2-medium models (in Section˜C.1) on one of our selected datasets can be performed efficiently on a single NVIDIA V100 or GTX1080TI-12GB GPU within a few hours. Additionally, computing the embeddings for a single layer of these models requires only forward passes, which takes approximately 10 minutes on the same hardware.
@@ 4.6 Limitations
While versatile, our framework has several limitations. One drawback is its comparatively high computational complexity; to get a reliable dimension estimate, we need to build a sufficiently large sub-sample of the latent space by performing forward passes on the input data and subsequently computing the neighborhoods and estimates. Being independent of the model training procedure, these computations can be performed in parallel to any potential fine-tuning, thus not slowing down the training process. Further work could focus on making the local estimate computation more efficient and developing local, stable, and computationally efficient descriptors.
An intrinsic limitation arises from the assumptions of our dimension estimates. The TwoNN estimate, which we apply locally, is known to accurately reflect the true local dimension only under strong assumptions on the local point sampling distribution: The point sample should come from a Poisson point process on a subspace of uniform dimension. Note that these assumptions and locally constant density for the hidden states of transformer models have been empirically demonstrated (Valeriani et al.,, 2023; Viswanathan et al.,, 2025). While the TwoNN method is non-parametric, the precise absolute value of the mean TwoNN estimate depends on the hyperparameter choices in selecting the latent space sub-sample and neighborhood size. This implies that our mean of local dimensions is of a “relative” nature, because we cannot directly compare the dimensions’ values between different model architectures. However, this is only of secondary importance for our method, as we are more interested in changes of dimension than in absolute values. Moreover, our framework is general, permitting the use of other estimation methods. Our working hypothesis is that high-dimensional neighborhoods reflect over-parameterized and poorly localized behavior, whereas very low LID indicates over-compression or memorization. Between these extremes lies a “sweet spot” in LID, where representations retain the essential degrees of freedom required for effective generalization. Empirically, this intermediate range correlates with peak generalization performance. While this correlation supports the interpretability of our analysis, establishing a causal link between LID reduction and improved generalization remains an open question for future work.
@@ 5 Conclusion
We introduce a novel geometric perspective on LLM training dynamics by measuring the local intrinsic dimension of contextual token embeddings. While the latent spaces of contextual language models exhibit regional variation in dimensionality, the mean local dimension provides a stable, interpretable summary across dataset splits. Across diverse tasks, a sustained drop in mean local dimension reliably suggests improved generalization, offering a practical, unsupervised diagnostic signal that, among other things, enables the detection of grokking. Such a marked reduction in intrinsic dimension in supervised downstream tasks highlights their strong compressive effect on the latent space. On the other hand, rising dimensions can point to a tradeoff between generalization and specialization, typically observed during overfitting. Our approach enables monitoring without supervision by labels, which is particularly valuable in low-resource settings where validation labels are limited or unavailable. Thus, our findings highlight the utility of geometric descriptors for monitoring and interpreting LLM behavior beyond traditional label-based metrics.
@@
To our knowledge, we are the first to investigate local dimensional shifts and their implications in training and fine-tuning contextual language models for different downstream tasks. The findings of this work enhance the ongoing discussion regarding the interpretability, adaptability, and generalizability of LLMs. Our work provides a foundation for designing better model architectures and developing interventions that utilize the insight that lower intrinsic dimensions benefit machine learning problems. Our measures could inform parameter choices such as LoRA-ranks in the spirit of (Ed-dib et al.,, 2024), and are broadly applicable for other (post)-training phases, for example, those which involve RL-tuning. While our current dimension estimation method is not differentiable (due to nearest-neighbor graph construction and the subsequent TwoNN estimator), we see designing a differentiable surrogate or proxy loss to encourage local compression during training as an exciting opportunity for future work. The locality of the dimension estimates promises further applications beyond the scope of the current work. One immediate avenue for investigation is to see to what extent the dimension can be used to detect which data the model has been trained on. Another important direction would be to find connections between the underlying meaning that the tokens carry and their corresponding dimension estimates.
@@ Appendix A Sensitivity Analysis
  • In Section˜A.1, we show that with increasing sample and neighborhood sizes, the TwoNN estimates tend to converge in our setting, and already a small portion of samples suffices for reasonable reliability.
  • In Section˜A.2 we show that the mean local estimates are stable under applying noise to the embedding vectors.
  • In Section˜A.3, we show that for masked language models, both masked and regular token embeddings lead to similar mean estimates.
All embeddings in the following analyses are computed using the regular token embeddings from the last layer of the RoBERTa-base model.
@@ A.1 Dependence on Hyperparameter Choices
Note that in our sub-sampling, we shuffle the dataset split using a given seed, and then truncate to a beginning segment of the specified size. Thus, when comparing different data sequence sub-sampling sizes for a fixed seed, the smaller sequence sub-samples are subsets of the larger ones. In all the experiments in this section, the size of token sub-sample is set to $N=$ 60 00060\,000, and the local neighborhood size parameter $L=128$. Note that for the smallest sequence sub-sample size $M=$ $2000$, not enough non-padding tokens are necessarily available to reach a token sample size of $N=$ 60 00060\,000, but with more sequences, we get this common token space size.
We take sequence sub-samples of each given size with 5 different seeds and select a random token sub-sample for each of these, on which the local TwoNN estimates are based. For the results, see Figure˜5 and Figure˜6.
For a given sequence sample size, the average of the mean local estimates, calculated across different sampling seeds, remains consistent between the various dataset splits of the same dataset. This observation indicates that the sampling process is robust to differences in how the data is divided. Furthermore, the estimates stabilize as the sequence sample size increases, suggesting convergence of the local mean estimates with larger samples. Based on these observations, we set $M$ to the values $7000$ or 10 00010\,000 for the natural language-based tasks in Sections˜4.1, 4.3 and 4.4, because it represents a practical compromise between computation effort and stability of the value at that scale. This value provides sufficient sampling density to capture the underlying data distribution comparably across datasets while maintaining computational efficiency.
@@
Figure 5: Boxplots of the mean local TwoNN estimates for different sequence sample sizes $M$ ranging from $2000$ to 16 00016\,000. We compute the mean estimates for 5 different sequence sub-sampling seeds for the three splits of the MultiWOZ dataset. Here and in all subsequent boxplots, the average of the mean estimates is depicted as green triangles, their median is the orange line; outliers are shown as circles.
@@
Here, we want to investigate the influence of the token sub-sample size on the local TwoNN estimate. We conduct our analysis on the MultiWOZ dataset. Since here we are interested in the effect of the second sub-sampling process, we want to fix the first sampling step for the sequences: We take the sub-sample of the first $M=$ 10 00010\,000 sequences from the MultiWOZ training, validation, and test split. From the resulting token embedding spaces, we then sample 100 000100\,000 points (ignoring padding and end-of-sequence (EOS) tokens). Finally, we compute the local estimates based on the first $N=$ 10 00010\,000, 20 00020\,000, …, 90 00090\,000, 100 000100\,000 points in the given sub-sample. Note that we are left with slightly fewer points in the local estimation step for the largest sub-sample size because of our de-duplication step in Algorithm 1. This means that in a given token sub-sample, the beginning segment of the first 10 00010\,000 points is the same even when increasing the sample size, and we compare the mean of the local estimates of this beginning segment.
The analysis of the datasets reveals a clear trend in the behavior of the truncated mean values and their associated standard deviations as the sample size increases, see Figure˜7. Here, the mean stabilizes quickly, and the standard deviation is comparatively small. This justifies our choice of N=60  [^0] N=\$60\,000\$ in the natural language dataset settings.
(a) MultiWOZ Training split
(b) MultiWOZ Validation split
(c) MultiWOZ Test split
@@
Having fixed sample sizes $M$ and $N$, Figure˜8 shows the distribution of TwoNN dimension estimates across varying locality scale parameters, represented by the number of neighbors $L$. A key observation is the relative stability of the estimates as the number of neighbors increases, with smaller locality scales (fewer neighbors) yielding more diverse estimate distributions. This indicates that local properties of the data significantly influence dimension estimates.
@@
(a) MultiWOZ Validation split
(b) Reddit Validation split
Figure 8: Distribution of the local token-wise TwoNN estimates over different locality scales $L$ (number of neighbors). The mean and quartiles of the resulting measure varying over the different tokens in a fixed subsample are shown here. The more global the estimates, the smaller the standard deviation in the resulting distribution.